Dennis Wingo: The project is moving forward. We have found a source (former Ampex Field Engineering Head) of further documentation for the drives. We are working out how to get this information to us this week. I have also added Charlie Byrne to our mailing list here as he is the original progenitor of our entire project with his 1965 memo regarding the utility of the analog data. Charlie has graciously offered to help and has a funded LASER contract relative to the film.
When we get images we will be sharing them with him for analysis. We have almost completed replacing the power supply capacitors (except for a shipment that got sent to my office in Alabama rather than to the lab in CA) and we have found procedures whereby we have calibrated the power supplies. Pictures of the machine with updated parts to follow.
We have a quote and timeline from VMI on the heads and are anxiously awaiting any word on the funding for the heads. We have sent the pinch rollers ( a mechanical drive component for the Capstan drive made of rubber) to a refurbishment house and will have a quote for that this week and have them back in about two or three weeks.
The folks at JPL have been incredibly helpful as well and have set up a private website for us that has everything that we need for our submittals to the Planetary Data System. Thanks to Sue Lavoie and the folks down there for being incredibly responsive.
We have our known good head and we have found two others that are in good enough shape to use for testing of the machine.
Kenneth Williams figured out a great way to mount smaller new caps where the big caps went.
We are also beginning debug and alignment procedures for the drive’s subsystems this week.
I am in Alabama this week and will be back on the evening of the 11th.
Dennis Wingo: I am doing another one today because a lot of things happened. First of all the students and Lynn Harper’s kids came in today and got 90% of all of the tapes out of the boxes and properly categorized. This is a major milestone that Keith enabled. Some of the students have volunteered to come in on Thursday to finish the job. After Pizza everyone was happy that they helped today.
We also received a lot of capacitors today and replaced a lot of parts in the machines. Every part that we replace and update like this makes the machine more stable in operations and function. Following will be pictures of the 90 volt Capstan power supply before and after the new caps. The newer ones are a pretty blue color.
Third, Ken Zin reached another technical milestone tonight by fixing a connector problem that we had that shorted out several signals. After one broken pin that was routed around with coax, the machine powered up and the servos are now locking into synchronization with their control signals, a major milestone to get a real tape to play. A picture of the monitor and the scope displaying the multiburst signal off of the test tape is shown tonight.
We expect a lot more parts tomorrow and Thursday that will be installed in power supplies and starting to replace them on some of the circuit cards. We have a long way to go, but we have also come a very long way. The students will be active tomorrow replacing parts in the machine as well as with a set of spares for the machine. Having known good working spare parts is critical to keeping downtime to a minimum once tape operations are underway on a daily basis. The machines are designed for rapid servicing and all of the students are learning the ins and outs of the machine.
We lament that we will be losing James as he, along with Austin and Kenneth, are doing a great job. I also want them to refurbish the spare vacuum pump and other key parts that will be used as spares and for possible components for the possible second machine. Ken is pretty confident that if the funding is made available that we can get a second drive up and running which could almost cut in half the time for doing the digitization of the images.
We have been hard at it for almost a month now. Ken is taking two days to go to Yosemite and I will be going home for one week next week. I will be back on the evening of the 11th.
Dennis Wingo: The project is rolling on. We have had some good progress on many fronts. Today we had a telecon with JPL (Rafael Alanis and Steven Adams) about the metadata tagging for the images produced from this effort. It was a good telecon and we expect to receive the templates and other information that will allow us to put together the images in the format that will properly integrate with the Planetary Data System.
We also had a major breakthrough today from a friend of a friend of a friends friend for finding documentation for the Ampex FR-900 that is our best drive that we have been working on and operating for the last couple of weeks. A former field service manager from Ampex was found who has the critical schematics and assembly drawings stored on Aperture cards (microfilm on a Hollerith punch card). We have sent him our assembly numbers from our equipment and he is getting a quotation for printing and digitizing the data. We will also have the one document from the Stanford Ampex archive by early next week. Therefore we are reasonably confident that we will have sufficient documentation to fully return to operation the primary FR-900A drive that we have been working on.
Today though, since Ken is troubleshooting individual circuits on the drive and this is not amenable to three or four people working on that simultaneously, I have tasked the students to shift around working with Ken and to completely disassemble and clean the second best drive, an Ampex 901 (that we now know was designed and built especially for the USAF and Eglin AFB) drive that we already have all the documentation on. We have repaired this drive with the top frame from another drive (pictures enclosed) and have washed the drives power supplies and other subsystems. We will attempt to return it to minimal operational status this week as we have time, while focusing on the primary FR-900A drive.
If the follow on funding is obtained to refurbish the heads and the operations, it is our hope that we might possibly have two operational drives (NO PROMISES) which would shorten the time to digitize the images. We will know within a week or two whether this path makes any sense.
Troubleshooting is beginning in ernest now with Ken tracing out circuits. He has found a couple of problems that have been fixed and more that are in process. This is just a very detailed, circuit by circuit process. We have ordered several thousand dollars worth of parts in the last few days and some have already been received and installed. Note in the latest pictures the blue capacitors that are replacing the silver ones. The newer ones are generally much smaller for the same capacitance so some machining of sleeves to make up the difference will be required. Pictures follow.
Tomorrow is tape inventory day with Keith Cowing leading the inventory team of students to get an accurate count of the tapes and then to sort and properly label all the tapes so that we can find the tape that we want when we want during the digitizing process.
All in all we are making steady progress and have reached all of our month one milestones and are well on the way with month two work.
Dennis Wingo: It has been a very busy several days, especially the last two days with several visits by many people small and great. We have still managed to get some work done and we have a working head (not the good one but a least one that works, and we have been able to read RF off of the heads. The servos on the machine are having a hard time syncing up properly, which we knew was going to happen due to their advanced age. We are evaluating whether or not the servos from the drive next to the working one are the same configuration and if so (we think that they are), we will take them apart and refurbish them first. This allows us to use the working machine to continue to debug the electronics, speeding the overall effort.
In the next day or so we will be ordering several thousand dollars worth of detail parts that we have been cataloging from the electronics bays. Today we received enough large electrolytic capacitors that we were able to replace the filter caps in all of the card racks and then reinstall them, allowing us to bring the entire system back together. Basically this allowed us today to test the entire system’s power supplies (six big power supplies). We now have at least a full compliment of working power supplies and two known good spare power supplies. Last week we were able to replace almost all of the relays in the system so little by little, things are coming back together.
We have been able to run tape and with a set of heads that we found (worn but still serviceable) we have been running a test tape to begin the process of debugging the signal system.
Our team is now together with Keith leading the effort to properly catalog and sort the tapes for our initial inventory. We are peeling off one of our students to help him and some more students volunteered to help with the work today (at least four)
In another coup today, one of the U.S.G.S. scientists from Flagstaff came by today (Dr. Lisa Gaddis), and came away a believer. Lisa has been working with the film and they have been doing a great job and they have offered help us with the selenographic details of properly registering the images. Couple this with the help from JPL to metatag the data for the Planetary Data System and we have the critical elements that we need to properly integrate the images into the mainstream of NASA image databases.
On top of all of this Charlie Byrne, the scientist from Bellcom that wrote the original justification for NASA to use analog tape drives to preserve a higher quality dataset for Lunar Orbiter came by today. We have his original memo where he laid out the justification for the tapes and we printed this and got him to sign it. It will be preserved and presented to the right people a the right time as a gift. Charlie also signed the drive today! That was an amazing thing and we put his photo on our facebook page today.
Additionally, the Google Moon folks came by yesterday and are seriously considering at least some level of support. Eric Anderson and Peter Diamandis were also by today.
Dennis Wingo: Ok, we have been sandbagging everyone just a little. We have been fixing and cleaning, and testing with power some of the subassemblies for the drive. We have also started reassembling the drive and for the first time since 1992 we have powered the drive up!!
Pictures to follow. However, everything is not all wonderful. We have started ordering parts and for some of them the lead time is weeks. This is going to slow down progress a lot soon. Also, there is a possibility that the drive that we have been working on may not be the one that we eventually return to working order.
The problem is that the documentation that we have is for the two older drives, not the much newer one that is in much better physical condition. This will cause a lot of problems in troubleshooting the newer drive if we can’t find the documents for the newer drive.
The drives originally came from Eglin AFB and we may try and see if they still have any documents or materials down there. We tried to call the contact person that we had from Eglin but he died last year. Ken Zin also called one of his contacts, who is in a nursing home and unable to help. We have called Ampex and tomorrow we will talk to the folks at the Stanford Ampex Museum.
There is a very large collection there and we may have to go dig in a warehouse to find the docs but it is a good possibility so all is not lost. This is how these projects go, part detective and part scrounging.
We are to the point where we can test some subsystems and will be doing more to do that over the next week or so.